# revtex4-2Repair the float

### v1

@@ I Introduction
The unprecedented science accessible through these experiments relies upon accurate measurements of the binary parameters, which, in turn, depend on the waveform models used to analyze the GW signals. Mismatched waveforms can potentially cause wrong interpretations of binary events and introduce biases in parameter estimation [16]. For instance, imperfect waveform models, based on semi-analytic approximations of two-body solutions in general relativity calibrated to numerical relativity simulations, can lead to systematic biases in the inferred binary parameters. Statistical errors decrease with signal-to-noise ratio while systematic errors do not, and therefore this issue will become more pressing as the sensitivity of GW detectors increases [17, 18, 19, 20, 21, 22].
@@
In this work we discuss the relevance of environmental effects in the interpretation of GW events and the mismatch with waveform models assuming that the binary evolves in vacuum, i.e., in the absence of an external environment. The effect of environments on GW signals driven by secular phenomena such as baryonic accretion, superradiance, and dark matter physics has been studied in Refs. [23, 24, 25, 26, 27, 28, 29]. In this paper, we mostly focus on the systematic biases induced by ignoring environmental tidal effects in gravitational waveforms.
@@
Tidal effects are widely recognized as a crucial factor in the inspiral of binary systems. Their leading contribution to the gravitational waveform appears at fifth post-Newtonian (PN) order [30, 31], and their measurement provides valuable insight into the internal structure of compact objects. For example, tidal effects have been used to probe the equation of state of neutron stars [32, 33], and they could reveal the existence of exotic compact objects or new physics [34, 35, 36, 37, 38, 39, 40, 41]. Within the realm of linear response theory, tidal effects are usually parameterized in terms of a set of coefficients called tidal Love numbers (TLNs) [42, 43, 44, 45] which, for various families of asymptotically flat BHs in four-dimensional space-times, are found to vanish exactly [46, 44, 45, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67]. This property has also been verified at the nonlinear level in Refs. [68, 69, 70, 71, 72, 73]. The vanishing of the TLNs can however be violated in scenarios involving BH mimickers [74, 34, 75, 76, 77, 78], in the presence of a cosmological constant, in theories of modified gravity, and in higher dimensions [79, 34, 80, 81, 82, 83, 84, 55, 85, 86, 87, 88, 89]. The nature of TLNs has also been investigated in the presence of time-dependent tidal fields [90, 91, 92, 93, 94, 95, 96].
@@
The manuscript is organized as follows. In Sec. II we provide the general framework to describe environmental effects in gravitational waveforms and outline the data analysis formalism. In Sec. III we compare waveforms in vacuum and in different environments for binaries detectable by ET and by LISA. Section IV is devoted to the conclusions. Some technical details are discussed in three appendices. Unless specified otherwise, we adopt geometrical units ($G=c=1$).
@@ II Environmental effects: setup and data analysis
We consider equal-mass binary BH systems with component masses $m_{1}=m_{2}=m_{\text{\tiny{\rm BH}}}$. Each BH is embedded in an external environment of mass $m_{\rm env}$, characterized by a fundamental length scale $L$. This environment could represent, for instance, an accretion disk or a dark matter halo, with $L$ corresponding to the location of the maximum density of the distribution. We assume the mass of the environment to be small compared to $m_{\text{\tiny BH}}$, i.e., we consider physical scenarios in which the parameter $\epsilon\equiv m_{\rm env}/m_{\text{\tiny{\rm BH}}}\ll 1$ and neglect terms of order $\mathcal{O}(\epsilon^{2})$.
@@
where $d_{\text{\tiny{\rm Roche}}}=\gamma\,m_{\text{\tiny{\rm BH}}}\tilde{L}$ is the Roche radius of each dressed BH. The numerical parameter $\gamma$ takes values ranging from 1.26 for rigid bodies to 2.44 for fluid ones [112]. We will assume $\gamma=2.44$, as it is more accurate in realistic scenarios. For comparison, we also give the ISCO frequency of the binary
@@
which is always bigger than $f_{\text{\tiny cut}}$ in our scenario. In Appendix A we discuss the interaction of the binary system with the environment after the fading process.
@@ II.1 Waveform model and parameter estimation
where $\bm{\theta}$ is the set of intrinsic and extrinsic source parameters. The amplitude ${\cal A}(f;\bm{\theta})$ and point-particle phase $\psi_{\text{\tiny PP}}$ are the outputs of the chosen spin-aligned waveform, which also depends on the individual spin components aligned with the orbital angular momentum of the binary, $\chi_{1,z}$ and $\chi_{2,z}$. In our analysis, we will assume nonspinning binaries ($\chi_{1,z}=\chi_{2,z}=0$), although the aligned spin components are included as waveform parameters.
@@
To predict the measurement accuracy of the source parameters, we use the Fisher information matrix approach [122, 123, 124]. In this framework, for GW signals with a high signal-to-noise ratio (SNR), the posterior distribution of the model parameters approximates a multivariate Gaussian centered at their true values, $\bm{\hat{\theta}}$, with a covariance matrix given by ${\Sigma}_{ij}=(\Gamma_{ij})^{-1}$, where
@@ III.1 Case 1: vacuum vs vacuum + TLN
$\simeq\rho\,\sigma_{\theta_{i}}\psi_{\text{\rm tidal}}(f_{\text{\tiny min}})\,,$
@@
The oscillations in $\Delta\theta_{i}$ observed at large values of $\tilde{\Lambda}$, and common to all panels of Fig. 1, are due to the magnitude of the tidal phase in Eq. (13). For large values of the tidal deformability, the difference $h_{\rm true}-h_{\rm rec}\sim(e^{{\rm i}\psi_{\rm tidal}}-1)$ cannot be linearized around $\tilde{\Lambda}$, leading to an oscillatory behavior.
@@ III.2 Case 2: vacuum vs fading environment
Similar to case 1, our analysis shows that the ratio $\Delta\theta_{i}/\theta_{i}$ remains much smaller than one across the entire parameter space we consider for the chirp mass. However, it becomes significant for both the symmetric mass ratio and the effective spin. The bias is strongly dependent on $\epsilon$. We computed $\Delta\theta_{i}$ for different values of this parameter (not shown here) and found that systematic errors scale linearly with $\epsilon$. A note of caution is needed: while biases on ${\cal M}$ can be small compared to ${\cal M}$ itself, this may not be the case when we compare them with the statistical uncertainty. Because $\cal M$ is the best-constrained mass parameter in the inspiral, uncertainties for ET and LISA can be as low as $\sigma_{\cal M}/{\cal M}={\cal O}(10^{-5})$ and $\sigma_{\cal M}/{\cal M}={\cal O}(10^{-5}-10^{-6})$, respectively, for sources at $d_{L}=1\,{\rm Gpc}$. Biases can thus be 10 times larger than the statistical uncertainty, especially for low $\tilde{L}$.
@@ IV Conclusions
There are numerous avenues for future research. A key objective is to expand our analysis by incorporating more sophisticated models of environmental effects, ranging from accretion disks to superradiance clouds, in order to gain a deeper understanding of the induced systematic biases. In relation to this point, note that an external environment may also induce (small) corrections to the tidal heating dissipative contribution in the BH response, which enter the waveform at a lower PN order [144], and thus have to be taken into account for a more complete analysis. Additionally, generalizing our results through full Bayesian analyses would provide valuable insight into the reconstruction of model parameters in the presence of dense environments surrounding binary BH systems. Finally, we aim to conduct an in-depth comparison of environmental effects with modified gravity theories within the context of tidal searches. These extensions will be explored in future work.
@@ Acknowledgements.
We thank Daniel D’Orazio, Konstantinos Kritos, Toni Riotto and Jay Wadekar for interesting discussions. V.D.L. thanks the Johns Hopkins University for the warm hospitality during the realization of this project. V.D.L. is supported by funds provided by the Center for Particle Cosmology at the University of Pennsylvania. E.B., L.D.G. and F.I. are supported by NSF Grants No. PHY-2207502, AST-2307146, PHY-090003 and PHY-20043, by NASA Grant No. 21-ATP21-0010, by the John Templeton Foundation Grant 62840, and by the Simons Foundation. L.D.G. is supported by Simons Investigator Grant No. 144924. The work of F.I. is supported by a Miller Postdoctoral Fellowship. E.B. and A.M. acknowledge support from the ITA-USA Science and Technology Cooperation program, supported by the Ministry of Foreign Affairs of Italy (MAECI). A.M. acknowledges financial support from MUR PRIN Grants No. 2022-Z9X4XS and No. 2020KB33TP.
@@ Appendix A Interactions with surrounding matter after the fading
The tidal disruption of the BH dress during coalescence is expected to pollute the binary surroundings. One might argue that the environmental debris could influence the system’s evolution by interacting with the binary components. In this section, we demonstrate that this is not the case, and that the dynamics after tidal disruption is dominated by GW emission. We consider two primary mechanisms that could affect the BH evolution: matter accretion and drag due to dynamical friction.
@@
For the benchmark value $\epsilon=10^{-3}$ used in this work, this ratio is small, provided that $\tilde{L}\lesssim 20$.
To summarize, matter accretion and dynamical friction are not expected to affect the binary dynamics after the Roche frequency. For similar reasons, corrections arising from effects entering the waveform at negative PN orders, such as gas migration [148], do not significantly affect the GW emission after depletion, for the values of $\tilde{L}$ we considered.

### v2

@@ I Introduction
The unprecedented science accessible through these experiments relies upon accurate measurements of the binary parameters, which, in turn, depend on the waveform models used to analyze the GW signals. Mismatched waveforms can potentially cause wrong interpretations of binary events and introduce biases in parameter estimation [16]. For instance, imperfect waveform models, based on semi-analytic approximations of two-body solutions in general relativity calibrated to numerical relativity simulations, can lead to systematic biases in the inferred binary parameters. Statistical errors decrease with signal-to-noise ratio (SNR) while systematic errors do not, and therefore this issue will become more pressing as the sensitivity of GW detectors increases [17, 18, 19, 20, 21, 22].
@@
In this work we discuss the relevance of environmental effects in the interpretation of GW events and the mismatch with waveform models assuming that the binary evolves in vacuum, i.e., in the absence of an external environment. The effect of environments on GW signals driven by secular phenomena such as baryonic accretion, superradiance, and dark matter physics has been studied in Refs. [23, 24, 25, 26, 27, 28, 29, 30, 31]. In this paper, we mostly focus on the systematic biases induced by ignoring environmental tidal effects in gravitational waveforms.
@@
Tidal effects are widely recognized as a crucial factor in the inspiral of binary systems. Their leading contribution to the gravitational waveform appears at fifth post-Newtonian (PN) order [32, 33], and their measurement provides valuable insight into the internal structure of compact objects. For example, tidal effects have been used to probe the equation of state of neutron stars [34, 35], and they could reveal the existence of exotic compact objects or new physics [36, 37, 38, 39, 40, 41, 42, 43, 44]. Within the realm of linear response theory, tidal effects are usually parameterized in terms of a set of coefficients called tidal Love numbers (TLNs) [45, 46, 47, 48] which, for various families of asymptotically flat BHs in four-dimensional space-times, are found to vanish exactly [49, 47, 48, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70]. This property has also been studied at the nonlinear level [71, 72, 73], with results demonstrating that it remains valid at every perturbative order in general relativity [74, 75, 76]. The vanishing of the TLNs can, however, be violated in scenarios involving BH mimickers [77, 36, 78, 79, 80, 81], in the presence of a cosmological constant, in theories of modified gravity, and in higher dimensions [82, 36, 83, 84, 85, 86, 87, 58, 88, 89, 90, 91, 92, 93]. The nature of TLNs has also been investigated in the presence of time-dependent tidal fields [94, 95, 96, 97, 98, 99, 100, 101, 102].
Figure 1: Schematic illustration of the binary system under consideration, consisting of two BHs (black circles) embedded in external environments (orange halos). In the left panel, prior to reaching the Roche frequency $f_{\text{\tiny cut}}$, the binary undergoes tidal interactions characterized by a nonzero tidal Love number. After tidal disruption (right panel), the BHs evolve as isolated (“naked”) objects within a significantly more diffuse halo. Further details are provided in Appendix A.
@@
The manuscript is organized as follows. In Sec. II we provide the general framework to describe environmental effects in gravitational waveforms and outline the data analysis formalism. In Sec. III we compare waveforms in vacuum and in different environments for binaries detectable by ET and by LISA. Section IV is devoted to the conclusions. Some technical details are discussed in three Appendixes. Unless specified otherwise, we adopt geometrical units ($G=c=1$).
@@ II Environmental effects: setup and data analysis
In this work, we investigate the dynamics of equal-mass binary BH systems surrounded by small-scale, localized environments, initially assumed to be in stationary equilibrium around each BH. Such environments are astrophysically motivated and may arise in a variety of contexts, including accretion disks, superradiant clouds, or dense dark matter substructures bound to individual BHs. This setup is physically distinct from scenarios involving circumbinary structures – such as those studied in Ref. [113] – where the binary evolves within a common disk from the outset [114]: for systems evolving in circumbinary structures, the effects of the environment are not limited to the ones considered in this paper, but they include also accretion, dynamical friction and drag forces, even during the early stages of the inspiral.
We consider equal-mass binary BH systems with component masses $m_{1}=m_{2}=m_{\text{\tiny{\rm BH}}}$. Each BH is embedded in an external environment of mass $m_{\rm env}$, characterized by a fundamental length scale $L$ corresponding, for instance, to the location of maximum density of the matter distribution. We assume the mass of the environment to be small compared to $m_{\text{\tiny BH}}$, i.e., we consider physical scenarios in which the parameter $\epsilon\equiv m_{\rm env}/m_{\text{\tiny{\rm BH}}}\ll 1$, and neglect terms of order $\mathcal{O}(\epsilon^{2})$. A schematic picture of the setup under investigation is sketched in Fig. 1.
@@
where $d_{\text{\tiny{\rm Roche}}}=\gamma\,m_{\text{\tiny{\rm BH}}}\tilde{L}$ is the Roche radius of each dressed BH. The numerical parameter $\gamma$ takes values ranging from 1.26 for rigid bodies to 2.44 for fluid ones [120]. We will assume $\gamma=2.44$, as it is more accurate in realistic scenarios. For comparison, we also give the GW frequency at the ISCO of the binary
@@
which is always bigger than $f_{\text{\tiny cut}}$ in our scenario. In Appendix A we discuss the interaction of the binary system with the tidally destroyed environment following the fading process, showing that GW emission predominantly drives the system’s dynamics for the parameter range considered in this study. This is due to the fact that, by the time the environment has been largely depleted and a circumbinary structure has formed, the binary’s orbital frequency has increased sufficiently, rendering environmental effects – such as accretion or dynamical friction – negligible.
@@ II.1 Waveform model and parameter estimation
where $\bm{\theta}$ is the set of intrinsic and extrinsic source parameters. The amplitude ${\cal A}(f;\bm{\theta})$ and point-particle phase $\psi_{\text{\tiny PP}}$ are the outputs of the chosen spin-aligned waveform, which also depends on the individual spin components aligned with the orbital angular momentum of the binary, $\chi_{1,z}$ and $\chi_{2,z}$. In our analysis we will assume nonspinning binaries ($\chi_{1,z}=\chi_{2,z}=0$), although the aligned spin components are included as waveform parameters.
@@
To predict the measurement accuracy of the source parameters, we use the Fisher information matrix approach [130, 131, 132]. In this framework, for GW signals with a high SNR, the posterior distribution of the model parameters approximates a multivariate Gaussian centered at their true values, $\bm{\hat{\theta}}$, with a covariance matrix given by ${\Sigma}_{ij}=(\Gamma_{ij})^{-1}$, where
@@ III.1 Case 1: vacuum vs vacuum + TLN
$\simeq\rho\,\sigma_{i}\psi_{\text{\rm tidal}}(f_{\text{\tiny min}})\,,$
@@
The oscillations in $\Delta\theta_{i}$ observed at large values of $\tilde{\Lambda}$, and common to all panels of Fig. 2, are due to the magnitude of the tidal phase in Eq. (13). For large values of the tidal deformability, the difference $h_{\rm true}-h_{\rm rec}\sim(e^{{\rm i}\psi_{\rm tidal}}-1)$ cannot be linearized around $\tilde{\Lambda}$, leading to an oscillatory behavior. Let us emphasize that, in such cases, the template-signal mismatch becomes sufficiently large that a real detection would likely reveal substantial fitting issues in the data analysis, rather than merely minor systematic errors in parameter estimation, ultimately preventing any meaningful recovery of the source parameters. This highlights that the results obtained for large values of $\tilde{\Lambda}$ should be viewed with caution when interpreted as systematic biases. A similar comment will also apply when considering large environmental scales $\tilde{L}$ in Case 3.
@@ III.2 Case 2: vacuum vs fading environment
Similar to case 1, our analysis shows that the ratio $\Delta\theta_{i}/\theta_{i}$ remains much smaller than one across the entire parameter space we consider for the chirp mass. However, it becomes significant for both the symmetric mass ratio and the effective spin. The bias is strongly dependent on $\epsilon$. We computed $\Delta\theta_{i}$ for different values of this parameter (not shown here) and found that systematic errors scale linearly with $\epsilon$. A note of caution is needed: while biases on ${\cal M}$ can be small compared to ${\cal M}$ itself, this may not be the case when we compare them with the statistical uncertainty. Because $\cal M$ is the best-constrained mass parameter in the inspiral, uncertainties for ET and LISA can be as low as $\sigma_{\cal M}/{\cal M}={\cal O}(10^{-5})$ and $\sigma_{\cal M}/{\cal M}={\cal O}(10^{-5}-10^{-6})$, respectively, for sources at $d_{L}=1\,{\rm Gpc}$. Biases can thus be 10 times larger than the statistical uncertainty, especially for low $\tilde{L}$. We emphasize that, for values of $\tilde{L}\lesssim\mathcal{O}(5)$, the environment would neither be in equilibrium nor gravitationally bound to the BHs, as discussed in Sec. II. In other words, in this regime, the component BHs would not be embedded in an external environment. While our analysis is restricted to $\tilde{L}\gtrsim\mathcal{O}(5)$ – where the presence of an external environment leads to an augmented mass and nonvanishing TLNs, and hence a nonzero bias – we do not model the environmental physics in detail at smaller scales $\tilde{L}$. In this limit, the inferred bias is generally expected to vanish, consistently with the absence of environmental effects (even though the very definition of bias may become ambiguous, as the system becomes fully degenerate with a BH binary of component masses $m_{\text{\tiny BH}}(1+\epsilon)$). This conclusion can be affected by the presence of strong correlations among the environmental and vacuum model parameters, which could possibly cause inference bias even when these effects are absent in the actual signal [147].
@@ IV Conclusions
There are numerous avenues for future research. A key objective is to expand our analysis by incorporating more sophisticated models of environmental effects, ranging from accretion disks to superradiance clouds, in order to gain a deeper understanding of the induced systematic biases. In relation to this point, note that an external environment may also induce (small) corrections to the tidal heating dissipative contribution in the BH response, which enter the waveform at a lower PN order [153], and thus have to be taken into account for a more complete analysis. Additionally, generalizing our results through full Bayesian analyses would provide valuable insight into the reconstruction of model parameters in the presence of dense environments surrounding binary BH systems, along the line of Refs. [27, 30]. Finally, we aim to conduct an in-depth comparison of environmental effects with modified gravity theories within the context of tidal searches. These extensions will be explored in future work.
@@ Acknowledgements.
We thank Daniel D’Orazio, Konstantinos Kritos, Toni Riotto and Jay Wadekar for interesting discussions. We also thank the anonymous referee for their valuable feedback on the manuscript. V.D.L. thanks the Johns Hopkins University for the warm hospitality during the realization of this project. V.D.L. is supported by funds provided by the Center for Particle Cosmology at the University of Pennsylvania. E.B., L.D.G. and F.I. are supported by NSF Grants No. PHY-2207502, AST-2307146, PHY-090003 and PHY-20043, by NASA Grant No. 21-ATP21-0010, by the John Templeton Foundation Grant 62840, and by the Simons Foundation. L.D.G. is supported by Simons Investigator Grant No. 144924. The work of F.I. is supported by a Miller Postdoctoral Fellowship. E.B. and A.M. acknowledge support from the ITA-USA Science and Technology Cooperation program, supported by the Ministry of Foreign Affairs of Italy (MAECI). A.M. acknowledges financial support from MUR PRIN Grants No. 2022-Z9X4XS and No. 2020KB33TP.
@@ Appendix A Interactions with surrounding matter after the fading
The tidal disruption of the BH dress during coalescence is expected to pollute the binary surroundings. One might argue that the environmental debris could influence the system’s evolution by interacting with the binary components, in a manner analogous to circumbinary structures (see the right panel of Fig. 1 for a schematic illustration). While a comprehensive understanding of such dynamics likely requires detailed simulations, here we aim to provide analytical insights demonstrating that the dynamics after tidal disruption is expected to be dominated by GW emission. Specifically, we consider several key mechanisms that could impact the evolution of the BHs – namely, matter accretion, dynamical friction, and drag forces – and show that the associated energy losses are subdominant compared to those from gravitational radiation.
@@
For the benchmark value $\epsilon=10^{-3}$ used in this work, this ratio is small, provided that $\tilde{L}\lesssim 20$. Furthermore, the energy loss by each BH due to pressure drag forces is roughly estimated as
$\dot{E}_{\text{\tiny{drag}}}\sim\rho_{\text{\tiny env}}\,m_{\text{\tiny{BH}}}^{2}v^{3}\sim\epsilon/\tilde{L}^{\nicefrac{{9}}{{2}}}\,.$ (28)
Hence, $\dot{E}_{\text{\tiny{drag}}}/\dot{E}_{\text{\tiny{GW}}}\sim\epsilon\,\tilde{L}^{\nicefrac{{1}}{{2}}}$. For $\epsilon=10^{-3}$ and $\tilde{L}\lesssim 20$, this ratio takes similar small values. To summarize, matter accretion, dynamical friction, and drag forces are not expected to affect the binary dynamics after the Roche frequency. For similar reasons, corrections arising from effects entering the waveform at negative PN orders, such as gas migration [157], do not significantly affect the GW emission after depletion, for the values of $\tilde{L}$ we considered.
Finally, we assess whether the assumption of zero initial spin remains valid throughout the binary evolution. While we have shown that accretion has a negligible effect on the BH masses, it could, in principle, cause significant spin-up of the binary components. However, we can demonstrate that this is not the case. Using the geodesic model for spin evolution [158, 159], the evolution of the BH spin $\chi=J/m_{\text{\tiny BH}}^{2}$ is given by
$\dot{\chi}=\left({\cal F}(\chi)-2\chi\right)\frac{\dot{m}_{\text{\tiny BH}}}{m_{\text{\tiny BH}}}\,$ (29)
in terms of ${\cal F}(\chi)\equiv L(m_{\text{\tiny BH}},J)/m_{\text{\tiny BH}}E(m_{\text{\tiny BH}},J)$, which is only a function of $\chi$ and depends on
$E(m_{\text{\tiny BH}},J)$ $=\sqrt{1-2\frac{m_{\text{\tiny BH}}}{3r_{\text{\tiny ISCO}}}}\,,$
$L(m_{\text{\tiny BH}},J)$ $=\frac{2m_{\text{\tiny BH}}}{3\sqrt{3}}\left(1+2\sqrt{3\frac{r_{\text{\tiny ISCO}}}{m_{\text{\tiny BH}}}-2}\right)\,.$ (30)
In these equations, the ISCO radius reads
$\frac{r_{\text{\tiny ISCO}}}{m_{\text{\tiny BH}}}=\left[3+Z_{2}-\sqrt{\left(3-Z_{1}\right)\left(3+Z_{1}+2Z_{2}\right)}\right]\,,$ (31)
with $Z_{1}=1+\left(1-\chi^{2}\right)^{\nicefrac{{1}}{{3}}}[\left(1+\chi\right)^{\nicefrac{{1}}{{3}}}+\left(1-\chi\right)^{\nicefrac{{1}}{{3}}}]$ and $Z_{2}=\sqrt{3\chi^{2}+Z_{1}^{2}}$
Assuming a constant accretion $\dot{m}_{\text{\tiny BH}}$, we can integrate the spin evolution equation (29) to get
$\int_{0}^{\chi_{\text{\tiny{env}}}}\frac{\mathrm{d}\chi}{{\cal F}(\chi)-2\chi}=\frac{\dot{m}_{\text{\tiny BH}}}{m_{\text{\tiny BH}}}\tau\,,$ (32)
where $\tau$ denotes the time to coalescence from the depletion of the environment, i.e. $\tau\sim 10^{-4}f_{\text{\tiny{cut}}}^{\nicefrac{{-8}}{{3}}}m_{\text{\tiny BH}}^{\nicefrac{{-5}}{{3}}}\sim 10^{-3}\,m_{\text{\tiny BH}}\,\tilde{L}^{4}$ [155], while $\chi_{\text{\tiny env}}$ indicates the final BH spin. In the small angular momentum limit (as a consequence of the mild mass growth), the integrand on the left-hand side can be simplified to
$\frac{1}{{\cal F}(\chi)-2\chi}\approx\frac{1}{108}\left(23\chi+12\sqrt{6}\right)\,,$ (33)
such that at leading order in $\chi_{\text{\tiny env}}$ one gets
$\chi_{\text{\tiny env}}\simeq 10^{-2}\left(\frac{\tilde{L}}{20}\right)^{\nicefrac{{5}}{{2}}}\left(\frac{\epsilon}{10^{-3}}\right)\,.$ (34)
This result shows that the spin growth is negligible for the range of parameters considered in this work, as expected from the negligible growth of the BH masses due to the suppressed accretion phase [160].